OpenStudy (anonymous):

(5n^2+7n^3)+(7n^3-9n^2)

OpenStudy (anonymous):

Jst an example Simplifying 8 + 3(-9n + -3) = -7(5n + -2) + -2 Reorder the terms: 8 + 3(-3 + -9n) = -7(5n + -2) + -2 8 + (-3 * 3 + -9n * 3) = -7(5n + -2) + -2 8 + (-9 + -27n) = -7(5n + -2) + -2 Combine like terms: 8 + -9 = -1 -1 + -27n = -7(5n + -2) + -2 Reorder the terms: -1 + -27n = -7(-2 + 5n) + -2 -1 + -27n = (-2 * -7 + 5n * -7) + -2 -1 + -27n = (14 + -35n) + -2 Reorder the terms: -1 + -27n = 14 + -2 + -35n Combine like terms: 14 + -2 = 12 -1 + -27n = 12 + -35n Solving -1 + -27n = 12 + -35n Solving for variable 'n'. Move all terms containing n to the left, all other terms to the right. Add '35n' to each side of the equation. -1 + -27n + 35n = 12 + -35n + 35n Combine like terms: -27n + 35n = 8n -1 + 8n = 12 + -35n + 35n Combine like terms: -35n + 35n = 0 -1 + 8n = 12 + 0 -1 + 8n = 12 Add '1' to each side of the equation. -1 + 1 + 8n = 12 + 1 Combine like terms: -1 + 1 = 0 0 + 8n = 12 + 1 8n = 12 + 1 Combine like terms: 12 + 1 = 13 8n = 13 Divide each side by '8'. n = 1.625 Simplifying n = 1.625

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Welcome